A bullet travelling at 500 m/s is brought to rest by an impulse of 50 N*s. What is the mass of the bullet?


1
Expert's answer
2021-06-15T15:37:40-0400

Impulse can be expressed as a change in momentum from one moment of time to another: "J = \Delta p = m v_2 - m v_1".

The initial speed is "v_1 = 500 \frac{m}{s}", and final is "v_2 = 0 \frac{m}{s}", therefore the mass of the bullet is: "m = \frac{J}{v_2} = \frac{50 N \cdot s}{500 \frac{m}{s}} = 0.1 kg = 100 g",